# Pattern formation in a coupled driven diffusive system

### v1

@@ Abstract
We investigate pattern formation in a driven mixture of two repulsive particles by introducing a Field-based Lattice Model (FLM), a hybrid model that combines aspects of the driven Widom-Rowlison lattice gas (DWRLG) and its statistical field theory [1, 2]. We find that the FLM effectively captures the bulk behavior of the DWRLG in both low- and high-density phases, suggesting that phase transitions in these models may share a universality class. Under the effect of the drive, the FLM additionally reveals an intermediate regime, not reported in the previous DWRLG studies, characterized by “irregular stripes” with widely fluctuating widths, contrasting with the “regular”, well-ordered stripes found at higher densities. In this intermediate phase, the system exhibits long-range order, predominantly perpendicular to the drive direction. To construct a continuum description, we derive two coupled partial differential equations via a gradient expansion of the FLM mean mass-transfer equations, supplemented with additive noise. Designing a numerical solver using the pseudospectral method with dealiasing and stochastic time differencing, we reproduce the low-density microemulsion phase (characterized by a non-zero characteristic wavenumber $q^{*}$) and perpendicular stripes at high density. We identify the non-zero difference in the characteristic velocities of the fields as a necessary condition for perpendicular stripe formation in the high-density phase. The continuum model also uncovers novel behaviors not previously observed in the FLM, such as stripes aligned parallel to the drive and chaotic patterns. This work highlights how the interplay of external drive, particle interactions, and noise can lead to a rich phenomenology in strongly driven binary mixtures.
@@ I Introduction
While statistical mechanics has excelled in describing systems at equilibrium, a significant portion of the real world operates far from equilibrium. Among the most fascinating properties of nonequilibrium systems are the complex spatiotemporal patterns they exhibit [3, 4, 5], including those associated with phase transitions and critical phenomena [6]. There has been growing interest in understanding how such patterns emerge and how they can be used to infer important properties of physical systems, whether active or driven externally.
@@
In the study of systems driven out of equilibrium, significant open questions remain, including how to define phase coexistence [7] or a thermodynamic quantity of the same usefulness as entropy at equilibrium [8]. In these driven systems, an external energy source, typically an applied driving force, maintains the system out of equilibrium. Well known examples include driven diffusive systems (DDS) which involve biasing the diffusion of one or more densities in various ways. Pioneering studies include simulations with one or two species (i.e., the 2-state Ising [9, 10, 11] or 3-state Potts [12] Blum-Emery-Griffiths [13] models) of particles placed on a lattice [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. DDS exhibit a range of intriguing behaviors [25, 26], including generic long-range correlations at all temperatures [14, 15] and critical properties governed by a nonequilibrium fixed point [27, 28]. At low temperatures, these models typically exhibit phase-separation into stripes which form along the drive direction [14, 15]. Recently, some of the authors have considered a driven version [1, 2] of the Widom-Rowlison lattice gas [29] (WRLG). The WRLG is a lattice gas model in which the two species of particles ($A$ and $B$, say) diffuse freely, apart from the constraint of strict exclusion of nearest neighbor (NN) AA-$B$ pairs. With no internal energy, the only control parameter of the equilibrium WRLG is the overall densities of the two species. Governed by entropy alone, this equilibrium system transitions, as the densities increase, from a homogeneous state of disorder to a one with phase separation, much like the Ising lattice gas [29]. When both species are driven in the same direction, striking features emerge [1, 2]: In the disordered phase, the AA-$B$ correlation function develops an intrinsic, preferred wavenumber ($\mathbf{q}^{*}\propto\hat{\mathbf{x}}$, the drive direction)! These correlations build up as the overall density increases, until long range order appears – in the form of alternating $A$- and $B$-rich stripes, characterized by $\mathbf{q}^{*}$. In stark contrast with the driven Ising case [15], these stripes are always perpendicular to the drive. Further, since $\mathbf{q}^{*}$ is independent of $L$, the system is not scale invariant.
@@
Notably, similar nonequilibrium phenomena, such as pattern formation and phase separation, are found in active matter, where scalar active mixtures have highlighted the importance of nonreciprocal couplings and characteristic length scales far from equilibrium [30, 31, 32, 33, 34]. Although the active matter literature is filled with interesting pattern formation examples, many of the observations already appear in the simpler externally driven systems. The driven WRLG (DWRLG) provides an especially interesting and simple setting for exploring patterns. The model was originally introduced and studied in detail in Ref. [1], where its phase behavior and stripe formation were characterized through simulations. More recently, a coarse-grained field-theoretic description of its disordered phase was formulated using the Doi-Peliti (DP) formalism [35, 36] in Ref. [2]. In the latter work, the field equations describe the nonequilibrium dynamics as two coupled driven diffusive systems: the sum and difference of $A$ and $B$ particle densities (referred to as the “density” and “charge” fields in the following), with different characteristic velocities. The field theory is able to capture key features observed in the simulations, including the development of a characteristic wavenumber $\mathbf{q}^{*}$. The work highlights the importance of noise in generating this characteristic length.
@@
The present study builds upon Refs. [1, 2], bridging the gap between them. More specifically, we seek a minimal continuum description, inspired by the DP formalism, of the ordered (heterogeneous) phase, where open questions remain regarding the structure of the equations of motion and whether they are capable of capturing the stripe orientation observed in lattice simulations. We also revisit the properties of the low-density phase. To this end, we first introduce a new Field-based Lattice Model (FLM) that preserves all the symmetries of the original DWRLG. We find excellent agreement with the phenomenology reported in Refs. [1, 2], strongly suggesting that both models belong to the same universality class. While discrete field theories (DFTs) are commonly employed in the quantum domain [37], stochastic DFTs (SDFTs) are far less explored; our lattice model provides a simple and straightforward example of an SDFT. Next, we perform a gradient expansion (GE) of the FLM master equation, leading to a pair of coupled field equations that, we shall argue, satisfactorily describe both low- and high-density regimes.
@@
The equations we derive are consistent with the DP and Martin-Siggia-Rose-Janssen-De Dominicis (MSRJD) field equations developed for the DWRLG in Ref. [2], but include additional terms necessary for describing a stable ordered phase. In the absence of an external drive, our equations of motion follow the structure of a two-component Cahn-Hilliard (CH)-like equation [38], but the presence of a drive explicitly breaks rotational symmetry, leading to stripe formation, or traveling waves, similar to phenomena observed in scalar active matter [30, 39]. We demonstrate here that aspects of the phenomenology reported in those works already manifest in a driven matter context, without requiring non-reciprocal particle interactions or particle activity.
@@
The remainder of the paper is organized as follows: In the next section, we introduce the FLM, providing a brief characterization of the model. In Sec. III, we derive the field equations via a GE, including explicit expressions for the linear coefficients, with additional details outlined in Appendix A. In Sec. IV, we numerically integrate these equations using the pseudospectral method combined with stochastic second-order Heun exponential time-differencing (PSSETD2H) [40, 41] and smooth dealiasing techniques [42], with further numerical details provided in Appendices B and C. Both deterministic and stochastic versions of the field equations are analyzed. One of the key numerical findings is that the difference in the characteristic velocity of the charge and density fields plays a crucial role in the nonequilibrium behavior of the model, supporting the conclusions of [2], but in the ordered phase as well. Finally, Sec. V presents our conclusions and outlines potential future directions.
@@ II Field-based Lattice Model
Early studies of Widom-Rowlinson systems were based on two distinct microscopic models, the original one of interacting particles in the continuum [29], and a lattice-gas model [43]. In more recent studies [1, 2], a stochastic field theoretic version was also considered, with the bridge between the two established either phenomenologically or using the Doi-Peliti formalism. All exhibit two distinct regimes: a disordered low-density phase and a high-density ordered phase. Here, we consider a hybrid model  [^1] The rudiments of this model were introduced in [1], where it played the role of an intermediary between simulations of the lattice gas and studies of a phenomenological field theory. In the present study, we investigate a full stochastic version of it., with continuous variables (densities, $\rho_{A,B}\in\left[0,1\right]$) defined on a $L\times L$ lattice  [^2] Generalizations to lattices in higher dimensions are straightforward.. This approach follows the spirit of coarse graining a lattice gas and will be referred to as the Field-based Lattice Model (FLM). Expecting it to be qualitatively similar to the WRLG, we perform simulations to explore its behavior. As we will show, the driven FLM appears to be richer than the DWRLG, in that a third regime seems to be present as the overall density is varied between the homogeneous and striped phases. In this regime, which we call the irregular stripe regime in the following, there is long-range order only in the direction perpendicular to the drive, with well-defined stripes, but of markedly different widths. In the next sections, we present a straightforward gradient expansion of the FLM to arrive at a field-theoretic description with equations of motion for the densities. Comparisons with those field theories studied earlier [1, 2] will be presented. We begin with a review of the driven WRLG.
@@ II.1 Model specification
The driven WRLG [1] consists of two species of particles, $A$ and $B$, diffusing on a square lattice of $L^{2}$ sites with periodic boundary conditions. An (integer valued) site $\mathbf{s}=\left(i,j\right)$ is either empty or occupied by a single particle, corresponding to the site-occupancy variable $n_{\mathbf{s}}$ taking values $0,+1$ or 1-1. In addition, the opposing species repel each other, implemented by the exclusion of nearest neighbor (NN) AA-$B$ pairs. In the stochastic dynamics, a random particle is chosen and moved to a randomly chosen NN or next NN (NNN) site, provided the destination site is empty and its NN sites are not occupied by a particle of the opposing species. When an external drive $\delta$ is imposed, particle moves in the +x+x (increasing $i$) direction are favored over moves in the x-x (decreasing $i$) direction. Following previous simulation studies, we set the numbers of $A$ and $B$ particles equal, so we have just two control parameters: the total density of particles, $\rho$, and the drive, $\delta$. For $\delta=0$ (the WRLG), the dynamics satisfies detailed balance and the properties of its stationary state can be predicted by considering entropy alone, as every allowed configuration is equally probable. Phase separation occurs when $\rho$ is increased beyond a critical density, $\rho_{c}^{\left(0\right)}\simeq 0.618(1)$, with critical behavior in the Ising universality class [29]. For $\delta>0$ (the DWRLG), detailed balance is violated, the stationary distribution is not easily accessible, while novel phenomena emerge in the nonequilibrium steady-state (NESS), for all densities (below, near, and above a $\delta$-dependent $\rho_{c}^{\left(\delta\right)}$, with mostly $\rho_{c}^{\left(\delta\right)}>\rho_{c}^{\left(0\right)}$) [1, 2].
The FLM differs mainly in having continuous (instead of discrete) variables at each site of a periodic $L\times L$ lattice, i.e., $\rho_{A,B}\in\left[0,1\right]$ instead of $n=0,\pm 1$. Both densities are conserved and we consider only the case of equal overall densities:
@@
    where $\Delta i=i^{\prime}-i$ and $\delta\in\left[0,1\right]$ is the drive parameter. In other words, hops along (opposite) the direction of the drive are enhanced (suppressed).
@@
    Once $\mathbf{s}^{\prime}$ is selected, select one of the species at random. In case $A$ is chosen, then transfer the following amount of $A$-density from $\mathbf{s}$ to $\mathbf{s}^{\prime}$:
@@
    where $\epsilon$ is a random variable  [^3] In our simulations, $\epsilon$ is chosen from a uniform distribution. From the results, $\epsilon$ serves only to redefine a diffusion length. Note that this multiplicative aspect of the noise may become relevant near the boundary values: $\rho_{A,B}=0,1$. in $\left[0,1\right]$, $\rho_{A}(\mathbf{s})$ is the density of $A$-particles at site $\mathbf{s}$, $h(\mathbf{s}^{\prime})=1-\rho_{A}(\mathbf{s}^{\prime})-\rho_{B}(\mathbf{s}^{\prime})$ is the vacancy density at the neighboring site $\mathbf{s}^{\prime}$, and $\not{\rho}_{B}(\mathbf{s}^{\prime})$ is defined as
@@
    where the product is over all nearest neighbors $\mathbf{s}^{\prime\prime}$ of $\mathbf{s}^{\prime}$, excluding the original site $\mathbf{s}$. Note that the product contains three factors for NN hops, and four in the case of hopping to a NNN. If instead species $B$ is chosen, the same rules apply, simply exchanging $A$ and $B$ in all expressions. The expression for $\Delta\rho_{A,B}(\mathbf{s})$ models AA-$B$ repulsion through an estimate of the probability that the nearest neighbors of the target site are free of $B$’s  [^4] The form of Eq. (3) and these suppression factors are motivated in part by the remarkably good agreement between a conjectured formula for the current density and the observed one in DWRLG: Eq. (8) and Fig. 5 in [1]..
These steps are iterated to generate the evolution of the system. We emphasize that $\Delta\rho_{A,B}(\mathbf{s})\leq h(\mathbf{s}^{\prime})$, so that it is impossible for a site to be “overfilled”. As usual, a Monte Carlo Step (MCS) is defined as $2L^{2}$ iterations, so that, on average, each species at each site is given one attempt to move some of its contents.

### v2

@@ Abstract
We investigate pattern formation in a driven mixture of two mutually repulsive particle species via a Field-based Lattice Model (FLM), a hybrid model that combines aspects of the driven Widom-Rowlison lattice gas (DWRLG) and its statistical field theory [1, 2]. We find that the FLM effectively captures the bulk behavior of the DWRLG in both low- and high-density phases, suggesting that phase transitions in these models may share a common universality class. Under the effect of a drive, the FLM additionally reveals an intermediate regime, not reported in previous DWRLG studies, characterized by irregular stripes with widely fluctuating widths, contrasting with the regular, well-ordered stripes found at higher densities. In this intermediate phase, the system exhibits long-range order predominantly perpendicular to the drive direction. To construct a continuum description, we derive two coupled partial differential equations for the particle densities via a gradient expansion of the FLM mean mass-transfer equations, supplemented with additive noise. Designing a numerical solver using the pseudospectral method with dealiasing and stochastic time differencing, we reproduce the low-density microemulsion phase (characterized by a nonzero characteristic wavenumber $q^{*}$) and stripes perpendicular to the drive at high density. We identify the nonzero difference in the characteristic velocities of the sum and difference of the particle densities as a necessary condition for perpendicular stripe formation in the high-density phase. The continuum model exhibits novel behaviors not observed in the FLM, such as stripes aligned parallel to the drive, and chaotic patterns. This work highlights how the interplay of external drive, particle interactions, and noise can lead to a rich phenomenology in strongly driven binary mixtures.
@@ I Introduction
While statistical mechanics has excelled in describing systems at equilibrium, a significant portion of the world operates far from equilibrium. Among the most fascinating properties of nonequilibrium systems are the complex spatiotemporal patterns they exhibit [3, 4, 5], including those associated with phase transitions and critical phenomena [6]. There has been growing interest in understanding how such patterns emerge and how they can be used to infer important properties of physical systems, whether active or driven externally.
@@
In the study of systems driven out of equilibrium, significant open questions remain, including how to define phase coexistence [7] or a thermodynamic quantity of the same usefulness as entropy at equilibrium [8]. A driven system is subject to an external force (or “drive”, for simplicity) that maintains it out of equilibrium. Examples include driven diffusive systems in which the drive biases particle diffusion along an axis with periodic boundaries, as in the pioneering studies of driven Ising, Potts, and Blume-Emery-Griffths models (in their lattice-gas representations) [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Driven diffusive systems exhibit a range of intriguing behaviors [20, 21], including generic long-range correlations at all temperatures [9, 10] and critical properties governed by a nonequilibrium fixed point [22, 23]. At low temperatures, these models typically exhibit phase-separation into stripes which form along the drive direction [9, 10]. Recently, some of the present authors have considered a driven version [1, 2] of the Widom-Rowlinson lattice gas [24] (WRLG). The WRLG is a lattice gas model in which the two species of particles ($A$ and $B$, say) diffuse freely, apart from the constraint of strict exclusion of nearest neighbor (NN) AA-$B$ pairs. With no interactions except volume exclusion, the only control parameters of the equilibrium WRLG are the overall densities of the two species. Governed by entropy alone, this equilibrium system transitions, as the densities increase, from a homogeneous disordered state to one with phase separation, much like the Ising lattice gas [24]. When both species are driven in the same direction, striking features emerge [1, 2]: In the disordered phase, the AA-$B$ correlation function develops an intrinsic, preferred wavenumber ($\mathbf{q}^{*}\propto\hat{\mathbf{x}}$, the drive direction)! Correlations grow as the overall density increases, until long-range order appears – in the form of alternating $A$- and $B$-rich stripes, characterized by $\mathbf{q}^{*}$. In stark contrast with the driven Ising case [10], these stripes are always perpendicular to the drive. Further, since $\mathbf{q}^{*}$ is independent of $L$, the system is not scale invariant.
@@
Notably, similar nonequilibrium phenomena, such as pattern formation and phase separation, are found in active matter, where scalar active mixtures have highlighted the importance of nonreciprocal couplings and characteristic length scales far from equilibrium [25, 26, 27, 28, 29]. Although the active matter literature is filled with interesting examples of pattern formation, many such phenomena already appear in simpler externally driven systems. The driven Widom-Rowlinson lattice gas (DWRLG) provides an especially interesting and simple setting for exploring patterns. The model was originally introduced and studied in detail in Ref. [1], where its phase behavior and stripe formation were characterized through simulations. More recently, a coarse-grained field-theoretic description of its disordered phase was formulated using the Doi-Peliti formalism [30, 31] in Ref. [2]. In the latter work, the field equations describe the nonequilibrium dynamics as two coupled driven diffusive systems: the sum and difference of $A$ and $B$ particle densities (referred to as the “density” and “charge” fields in the following), with different characteristic velocities. The field theory is capable of capturing key features observed in the simulations at low particle densities, including the development of a characteristic wavenumber $\mathbf{q}^{*}$, and further highlights the importance of noise in generating this characteristic wavenumber in the disordered phase.
@@
The present study builds upon Refs. [1, 2], bridging the gap between them. More specifically, we seek a minimal continuum description, inspired by the Doi-Peliti formalism, of the ordered (heterogeneous) phase, where open questions remain regarding the structure of the equations of motion and whether they are capable of capturing the stripe orientation observed in lattice simulations. We also revisit the properties of the low-density phase. To this end, we first introduce a new Field-based Lattice Model (FLM) that preserves all the symmetries of the original DWRLG. We find excellent agreement with the phenomenology reported in Refs. [1, 2], strongly suggesting that both models belong to the same universality class. While discrete field theories are commonly employed in the quantum domain [32], their stochastic counterparts are far less explored; our lattice model provides a simple and straightforward example of a stochastic discrete field theory. Next, we perform a gradient expansion (GE) of the FLM master equation, leading to a pair of coupled field equations that, we shall argue, satisfactorily describe both low- and high-density regimes.
@@
The equations we derive are consistent with the Doi-Peliti and Martin-Siggia-Rose-Janssen-De Dominicis (MSRJD) field equations developed for the DWRLG in Ref. [2], but include additional terms necessary for describing a stable ordered phase. In the absence of an external drive, our equations of motion follow the structure of a two-component Cahn-Hilliard-like equation [33]. But a drive breaks rotational symmetry, leading to stripe formation, or traveling waves, similar to phenomena observed in scalar active matter [25, 34]. We demonstrate here that aspects of the phenomenology reported in [25, 34] are already evident in a driven matter context, without requiring non-reciprocal particle interactions or particle activity.
@@
The remainder of the paper is organized as follows: In the next section, we introduce the FLM, providing a brief characterization of the model. In Sec. III, we derive the field equations via a GE, including explicit expressions for the linear coefficients, with additional details outlined in Appendices A and B. In Sec. IV, we integrate these equations numerically using the pseudospectral method combined with stochastic second-order Heun exponential time-differencing (PSSETD2H) [35, 36] and smooth dealiasing techniques [37], with further numerical details provided in Appendices C and D. Both deterministic and stochastic versions of the field equations are analyzed. One of the key numerical findings is that the difference in the characteristic velocity of the charge and density fields plays a crucial role in the nonequilibrium behavior of the model, supporting the conclusions of [2], but in the ordered phase as well. Finally, Sec. V presents our conclusions and outlines potential future directions.
@@ II Field-based Lattice Model
Early studies of Widom-Rowlinson systems were based on two distinct microscopic models, the original one of interacting particles in the continuum [24], and a lattice-gas model [38]. In more recent studies [1, 2], a stochastic field theoretic version was also considered, with the bridge between the two established either phenomenologically or using the Doi-Peliti formalism. All exhibit two distinct regimes: a disordered low-density phase and a high-density ordered phase. Here, we consider a hybrid model  [^1] The rudiments of this model were introduced in [1], where it played the role of an intermediary between simulations of the lattice gas and studies of a phenomenological field theory. In the present study, we investigate a full stochastic version of it., with continuous variables (densities, $\rho_{A,B}\in\left[0,1\right]$) defined on a $L\times L$ lattice  [^2] Generalizations to lattices in higher dimensions are straightforward.. This approach follows the spirit of coarse graining a lattice gas and will be referred to as the Field-based Lattice Model (FLM). Expecting it to be qualitatively similar to the WRLG, we perform simulations to explore its behavior. As we will show, the driven FLM has a richer phenomenology than the DWRLG, in that a third regime appears as the overall density is varied between the homogeneous and striped phases. In this regime, which we call the irregular stripe regime in the following, there is long-range order only in the direction perpendicular to the drive, with well-defined stripes, but of markedly different widths. In the next sections, we present a straightforward gradient expansion of the FLM to arrive at a field-theoretic description with equations of motion for the densities, and compare our results with the field theories studied earlier [1, 2]. We begin with a review of the driven WRLG.
@@ II.1 Model specification
The driven WRLG [1] consists of two species of particles, $A$ and $B$, diffusing on a square lattice of $L^{2}$ sites with periodic boundary conditions. Each site $\mathbf{s}=\left(i,j\right)$ is either empty or occupied by a single particle. In addition, opposite species repel each other: nearest neighbor (NN) AA-$B$ pairs are forbidden. In the stochastic dynamics, a random particle is chosen and moved to a randomly chosen NN or next NN (NNN) site, provided the destination site is empty and its NN sites are not occupied by a particle of the opposite species. When an external drive $\delta$ is imposed, particle moves in the +x+x (increasing $i$) direction are favored over moves in the x-x (decreasing $i$) direction. Following previous simulation studies, we set the numbers of $A$ and $B$ particles equal, so that we have only two control parameters: the total particle density, $\rho$, and the drive, $\delta$. For $\delta=0$ (the WRLG), the dynamics satisfies detailed balance and the properties of its stationary state can be predicted by considering entropy alone, as every allowed configuration is equally probable. Phase separation occurs when $\rho$ is increased beyond a critical density, $\rho_{c}^{\left(0\right)}\simeq 0.618(1)$, with critical behavior in the Ising universality class [24]. For $\delta>0$ (the DWRLG), detailed balance is violated, the stationary distribution is unknown a priori, while novel phenomena emerge in the nonequilibrium steady-state, for all densities (below, near, and above a $\delta$-dependent $\rho_{c}^{\left(\delta\right)}$, with $\rho_{c}^{\left(\delta\right)}>\rho_{c}^{\left(0\right)}$, in general) [1, 2].
The FLM differs from the DWRLG mainly in having continuous (instead of discrete) variables at each site of the lattice, i.e., $\rho_{A,B}\in\left[0,1\right]$ instead of $0,\pm 1$. Both densities are conserved; here we consider only the case of equal overall densities:
@@
    where $\Delta i=i^{\prime}-i$ and $\delta\in\left[0,1\right]$ is the drive. In other words, hops along (opposite) the direction of the drive are enhanced (suppressed).
@@
    Next, select one of the species at random. In case $A$ is chosen, then transfer the following amount of $A$-density from $\mathbf{s}$ to $\mathbf{s}^{\prime}$:
@@
    where $\epsilon$ is a random variable  [^3] In our simulations, $\epsilon$ is chosen from a uniform distribution. From the results, $\epsilon$ serves only to redefine a diffusion length. Note that this multiplicative aspect of the noise may become relevant near the boundary values: $\rho_{A,B}=0,1$. in $\left[0,1\right]$, $\rho_{A}(\mathbf{s})$ is the density of $A$-particles at site $\mathbf{s}$, $h(\mathbf{s}^{\prime})=1-\rho_{A}(\mathbf{s}^{\prime})-\rho_{B}(\mathbf{s}^{\prime})$ is the vacancy density at $\mathbf{s}^{\prime}$, and $\not{\rho}_{B}(\mathbf{s}^{\prime})$ is defined as
@@
    where the product is over all nearest neighbors $\mathbf{s}^{\prime\prime}$ of $\mathbf{s}^{\prime}$, excluding the original site $\mathbf{s}$. Note that this product contains three factors for NN hops, and four in the case of hopping to a NNN. The factor $\not{\rho}_{B}$ models AA-$B$ repulsion through an estimate of the probability that the nearest neighbors of the target site are free of $B$’s  [^4] The form of Eq. (3) and these suppression factors are motivated in part by the remarkably good agreement between a conjectured formula for the current density and the observed one in DWRLG: Eq. (8) and Fig. 5 in [1].. If instead of $A$, species $B$ is chosen, the same rules apply, simply exchanging $A$ and $B$ in all expressions.
These steps are iterated to generate the evolution. We emphasize that $\Delta\rho_{A,B}(\mathbf{s})\leq h(\mathbf{s}^{\prime})$, so that it is impossible for a site to be “overfilled”. As usual, a Monte Carlo Step (MCS) is defined as $2L^{2}$ iterations, so that, on average, each species at each site is given one attempt to transfer some of its contents.